OpenStudy (anonymous):

(2/7 r^5 s^3)^2 Simplify? I've tried problems similar to this but the website keeps marking it wrong

OpenStudy (whpalmer4):

is that \[(\frac{2}{7}r^5s^3)^2\]or something else?

OpenStudy (anonymous):

Yeah, I couldn't get it like that

OpenStudy (whpalmer4):

Okay, what have you tried as an answer?

OpenStudy (whpalmer4):

Let's see if we can identify where you are going astray...

OpenStudy (anonymous):

\[(\frac{ 2 }{ 7 }r^5s^3)(\frac{ 2 }{ 7 }r^5s^3)\]

OpenStudy (anonymous):

I got 4/49 r^25 s^9

OpenStudy (anonymous):

Astray?

OpenStudy (anonymous):

Ooh I see when you expanded you multiplied the variables

OpenStudy (anonymous):

The problem is your exponents. (a^n)^m=a^(nm). You were squaring the exponents themselves. 2/7 ^2=4/49. r^5 ^2=r^10. s^3 ^2=s^6. Your answer should be 4/49 r^10 s^6

OpenStudy (jdoe0001):

\(\bf \left(\cfrac{2}{7}r^5 s^3\right)^{\color{red}{ 2}}\implies \left(\cfrac{2}{7}\right)^{\color{red}{ 2}}\cdot r^{5\cdot {\color{red}{ 2}}}s^{3\cdot {\color{red}{ 2}}}\implies \cfrac{2^{\color{red}{ 2}}}{7^{\color{red}{ 2}}}\cdot r^{5\cdot {\color{red}{ 2}}}s^{3\cdot {\color{red}{ 2}}}\)

OpenStudy (anonymous):

Sorry lost connection, nice job @jdoe0001 :D

OpenStudy (whpalmer4):

Yeah, I got cut off as well. When multiplying exponentials with the same base, you add the exponents (and subtract for division). This is the same as with logarithms, where when multiplying you add the logarithms, and when dividing you subtract them. Remember that for when you learn about logarithms :-)